Let A = {a,b,c,d} and B = {c,d,e,f,g}.
Let R1 = {(a,c), (b,d), (c,e)}
R2 = {(a,c), (a,g), (b,d), (c,e), (d,f)}
R3 = {(a,c), (b,d), (c,e), (d,f)}
Justify which of the given relation is a function from A to B.
(c) Let f be a real valued function defined by f(x) = 1
x2−9
.
(i) What is the domain of f?
(ii) What is the range of f?
(iii) Represent f as a set of ordered pairs.
1
Expert's answer
2020-08-13T18:33:36-0400

Given "A = \{a,b,c,d\}" and "B = \{c,d,e,f,g\}" .

"R_1 = \{(a,c), (b,d), (c,e)\}, R_2 = \{(a,c), (a,g), (b,d), (c,e), (d,f)\}, \\ R_3 = \{(a,c), (b,d), (c,e), (d,f)\}"

A relation is a function when every element of set A has image in B and a element of set A can-not have more than one image in set B.

So, Relation "R_3" is a function.


(c) Given "f" is a real valued function defined by "f(x) = x^2 - 9" .

(I) Function is defined for all real values of "x". Hence,

Domain of"f = \R"

(ii) Now, as "x^2 \geq 0 \implies x^2-9 \geq -9"

Hence, Range of "f = [-9,\infin)"

(iii) Representation of "f" as a set of ordered pair = "\{ (x,x^2-9) : x \in \R\}"